OpenStudy (anonymous):

Solve the system. (I need a step by step) y=5x-2 y=x+6

OpenStudy (anonymous):

y=5x-2 substitute y in equation 2 5x-2=x+6 5x-x=6+2 4x=8 x=8/4 x=2 now plug x value into one of the equation y=2+6 y=8 so x=2 and y=8

OpenStudy (anonymous):

Solving by canceling terms: subtract the second equation from the first. Results in (y-y)=(5x-x)+(-2-6) -> 0=4x-8. Add 8 to both sides to get 4x=8, divide both sides by 4 to get x=2. Next, plug this back in to the second equation, so y=2+6=8. So the answer to the system is (2,8) or x=2, y=8.

OpenStudy (anonymous):

Agh thanks so much. Saved my life.